OpenStudy (anonymous):

Evaluate -2 + 4 x 3^2 - 6

OpenStudy (kxttxn):

remember order of operations: PEMDAS (parenthesis, exponent, multiplication/division, addition/subtraction) so start with the exponent (3^2) and go from there

OpenStudy (anonymous):

yeah i got -2 + 4 x 9 - 6= -2 + 36 - 6 = 38 - 6+ 32

OpenStudy (kxttxn):

when you get to -2+36-6 you added the 2 on, but since it is negative it would be subtracted

OpenStudy (anonymous):

so what would i do next?

OpenStudy (anonymous):

The answer is 28

OpenStudy (kxttxn):

so instead of 38-6, it would be 34-6

OpenStudy (anonymous):

okay, i got it, ty :))

OpenStudy (kxttxn):

no problem :)